Circular No. 8707

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006bq, 2006bu, 2006by, 2006bz, 2006ca Three additional supernovae have been found in LOSS/KAIT images

(cf. IAUC 8705) by E. Lee, R. Chornock, J. Burket, M. Baek, and W. Li:

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006by May 3.31 13^{h}26^{m}09.^{s}12 +35^{o}55′58′′.0 18.2 0′′.1 E, 5′′.7 S
2006bz May 4.30 13 00 43.38 +27 57 41.8 17.5 7′′.0 E, 5′′.9 S
2006ca May 4.49 18 22 54.15 +12 26 03.2 16.4 14′′.0 E, 17′′.3 N

Additional magnitudes for 2006by in NGC 5149: Feb. 9.50 UT, [19.0; May 4.31, 18.2. Additional magnitudes for 2006bz in PGC 44809 (in Abell Cluster 1656): Jan. 13.46, [18.5; Feb. 5.43, [18.0; May 5.37, 17.1. Additional magnitudes for 2006ca in UGC 11214: 2005 Oct. 5.19, [19.0; 2006 Apr. 14.53, [17.0; May 5.40, 16.1. Additional KAIT magnitudes for 2006bq (cf. IAUC 8705): 2005 Oct. 5.20, [19.5; 2006 May 4.52, 16.8; 5.40, 17.0.

R. J. Foley, J. M. Silverman, M. Moore, and A. V. Filippenko, University of California, Berkeley, report that inspection of CCD spectra (range 330–1040 nm), obtained on May 5 UT with the Lick 3-m telescope (+ Kast), shows that SNe 2006bq and 2006bu (cf. IAUC 8705) are of type Ia at ∼ 20 and 0 days past maximum brightness, respectively. Spectra of SNe 2006by and 2006ca reveal they are both of type II. The spectrum of SN 2006ca shows a very blue continuum with weak Balmer absorption lines, indicating that it is a young object. After removing the recession velocity of 5680 and 2670 km/s determined from narrow emission lines, the minima of the Hα P-Cyg absorption is seen to have velocities of 7600 and 17200 km/s for SNe 2006by and 2006ca, respectively. A spectrum of SN 2006bz shows it to be of type Ia, resembling SN 1991bg (Filippenko et al. 1992, A.J. 104, 1543) at maximum light; removing the host-galaxy recession velocity of 8366 km/s (Caldwell et al. 1992, A.J. 106, 473), the minimum of the Si II 6150 absorption is measured to be 10800 km/s.

S. Blondin, M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithson-ian Center for Astrophysics, report that a spectrum (range 350–740 nm) of SN 2006bu, obtained on May 5.19 UT by W. Peters with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-Ia supernova, with a spectral-feature age (Riess et al. 1997, A.J. 114, 722) of 2 days past maximum light. Cross-correlation with a library of type-Ia supernova spectra indicates that SN 2006bw is at a redshift of cz ≈ 25000 km/s. Adopting this recession velocity for the supernova, the maximum absorption in the Si II line (rest 635.5 nm) is blueshifted by 11500 km/s.

2006 May 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green